OpenStudy (anonymous):

What is the equation for a line that passes through the points (-4, -4) and (8, 23)?

OpenStudy (er.mohd.amir):

first find slope m=(y2-y1)/(x2-x1) then eq is (y-y1)=m(x-x1)